OpenStudy (anonymous):

let T be multiplication by the matrix A= (top: 1,0 -2; bottom: -2,1,2) a) find a basis for the range of T b) find a basis for the kernel of T c) the rank and nullity of T d) the rank and nullity of A Please, explain me.

OpenStudy (anonymous):

For the basis of the range of \(T\), just look at \(T\begin{bmatrix}x\\y\\z\end{bmatrix}\):$$T\begin{bmatrix}x\\y\\z\end{bmatrix}=\begin{bmatrix}1&0&-2\\-2&1&2\end{bmatrix}\begin{bmatrix}x\\y\\z\end{bmatrix}=\begin{bmatrix}x+0y-2z\\-2x+1y+2z\end{bmatrix}=\begin{bmatrix}1\\-2\end{bmatrix}x+\begin{bmatrix}0\\1\end{bmatrix}y+\begin{bmatrix}-2\\2\end{bmatrix}z$$which means every vector in our range is a linear combination of our column vectors, i.e. the range is spanned by our column vectors; in fact, the column space of our matrix is exactly the range of our linear transform. Since we know \(\begin{bmatrix}1\\-2\end{bmatrix},\begin{bmatrix}0\\1\end{bmatrix},\begin{bmatrix}-2\\2\end{bmatrix}\) span our range, we can use them to determine a basis by picking linearly independent vectors from them. Since the collection of all three is dependent, we kick out \(\begin{bmatrix}-2\\2\end{bmatrix}\) and keep \(\begin{bmatrix}1\\-2\end{bmatrix},\begin{bmatrix}0\\1\end{bmatrix}\) as our basis since they're linearly independent.

OpenStudy (anonymous):

oops, my earlier line got cut off.$$\begin{bmatrix}x+0y-2z\\-2x+1y+2z\end{bmatrix}=\begin{bmatrix}1\\-2\end{bmatrix}x+\begin{bmatrix}0\\1\end{bmatrix}y+\begin{bmatrix}-2\\2\end{bmatrix}z$$

OpenStudy (anonymous):

I got that stuff but come up in another way, I take rref A to get (1,-2) and (0,1) . friend, I cannot "kick out " something without any reason or any logic. Anyway, i am there

OpenStudy (anonymous):

@Hoa we kick it out because we understand that the other two vectors are linearly independent without it.

OpenStudy (anonymous):

surely we understand, but my prof wants the steps to get this, so, the range of T is 2 right?

OpenStudy (anonymous):

The rank is the dimension of the range, which is indeed 2.

OpenStudy (anonymous):

The nullity is then the dimension of our kernel (null space). Together, by the rank-nullity theorem, they sum to the dimension of our linear transformation itself! :-) (which is \(3\)). This means the dimension of our kernel is \(3-2=1\).

OpenStudy (anonymous):

oh oh, we mess up there, the rref is \[\left[\begin{matrix}1 & 0&-2 \\ 0 & 1& -2\end{matrix}\right]\] so the basis for the range of T is \[\left(\begin{matrix}1 \\ 0\end{matrix}\right) and \left(\begin{matrix}0 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

So all we need is one vector in the null space of our matrix and we will have a basis. So, @Hoa, can you find one vector in the null space? :-) hint: reduce

OpenStudy (anonymous):

@Hoa both are valid bases for \(\mathbb{R}^2\).

OpenStudy (anonymous):

ok, x-2z =0 and y-2z =0

OpenStudy (anonymous):

Consider:$$\begin{bmatrix}1\\-2\end{bmatrix}+2\begin{bmatrix}0\\1\end{bmatrix}=\begin{bmatrix}1\\0\end{bmatrix}$$ ... so they definitely are a fine basis :-) they're linearly independent and span our vector space, which is all that is needed (though yours is also orthonormal and canonical)

OpenStudy (anonymous):

and z is free variable and I can express that free by \[\left(\begin{matrix}2 \\ 2\\1\end{matrix}\right)\]

OpenStudy (anonymous):

Exactly correct @Hoa :-) which is then a fine basis for our kernel as its dimension is \(1\).

OpenStudy (anonymous):

hi , what caniconal mean?

OpenStudy (anonymous):

standard

OpenStudy (anonymous):

ok, got it

OpenStudy (anonymous):

A linear transform and the matrix underlying the transform have the same rank/nullity.

OpenStudy (anonymous):

you mean, A and T?

OpenStudy (anonymous):

that is a specific example, yes

OpenStudy (anonymous):

so, in conclusion, with that kind of question, I have to get rref of A , consider the column space to get the pivots and free variable, number of pivots is the range of T , number of free is the ker (T) and base on that stuff, figure out the basis of them, right?

OpenStudy (anonymous):

Sure... but in linear algebra there are often many neat approaches to a problem.

OpenStudy (anonymous):

Thanks for your help. I appreciate. I rarely get help from that field.

OpenStudy (anonymous):

from calculus, many good friends there, but to linear, it is not hard but complicated, and it sounds like they don't want to spend time with that. I don't know why

OpenStudy (anonymous):

By the way, I typically try figuring out trivial linear combinations of the other column vectors I suspect of ruining linear independence... e.g. I merely did the following in my head:$$-2\begin{bmatrix}2\\-2\end{bmatrix}+2\begin{bmatrix}0\\1\end{bmatrix}=\begin{bmatrix}-4+2\\4+2\end{bmatrix}=\begin{bmatrix}-2\\2\end{bmatrix}$$

OpenStudy (anonymous):

nope, you make mistake there, right the first one -2*2 + 2 *0 = -4 , not -2

OpenStudy (anonymous):

ooops